# Beyond Single-Turn: A Survey on Multi-Turn Interactions with Large Language Models

### v4

@@ 2.2.4 Conversational Engagements in Jailbreak
Most prior research has focused on single-turn jailbreaks, in which adversaries use a single prompt—often few-shot—to trigger harmful content. For example, optimization-based methods in [166, 167] utilize gradient-derived gibberish suffixes to elicit such outputs, but these techniques depend on internal token probability knowledge and are not applicable to closed-source models. To explore alternative strategies, subsequent studies [168, 169] have investigated unconventional communication patterns, such as role-playing scenarios, and developed multi-turn conversational methods that leverage the entire dialogue history. These approaches demonstrate that multi-turn jailbreaking, where even seemingly innocuous prompts contribute to later interventions, presents a far more complex challenge.
@@
Multi-turn jailbreak can employ more diverse strategies. “Crescendo” [170, 171] is one of the strategies, which implicitly instruct victim LLMs to provide harmful information. The intuition behind this strategy is that LLMs agreeing to a small, initial request increases the likelihood of complying with subsequent, larger demands. By asking innocuous but implicitly suggestive questions, the token distribution of LLMs shifts towards the direction where tokens containing more harmful information are more likely to be generated. The multi-turn interactions as a whole jailbreaks LLMs instead of one specific questions plays the role. Besides, the implicitness of questions and the order to present the questions to LLMs are important to the success of jailbreak. Following the crescendo idea, Ren et al. (2024) [171] generalizes the crescendo method used in [170], where in [170] fixed and human-crafted seed instances are needed to generate attacks, making it challenging to generate diverse attacks. In [171], the novelty focuses on the process of self-discovering diverse attack clues inside the model’s prior knowledge via network structures and the classification of the clues, by constructing the two-layer relation tree according to Latour’s actor-network theory.
@@
Another important strategy for multi-turn jailbreaks is to decompose the one harmful prompt into several pieces where each contains less malicious contents [172, 173, 174]. Therefore, language models can incrementally generate harmful content through multi-turn dialogue. By decomposing the original malicious query into several less harmful sub-questions can evade the guardrail of LLMs and induce harmful responses. Due to the in-context learning capabilities, harmful knowledge can be gathered together in the final turn. Using such decomposition, alignment is quite successful for each turn in a multi-turn dialogue, except the final turn when all responses are gathered and combined. However, the cumulative harmful content across the dialogue results in an overall alignment failure. Following the idea, Wang et al. (2025) [175] trains a red-team jailbreak agent through the interactions with target LLMs to generate decomposed yet coherent jailbreak prompts. [176] proposes a scenario-based jailbreak method to disguise the malicious intent from guardrails of LLMs, assuming LLMs can only detect direct harmful intent but would be misled if the attacker creates a scenario claiming that others are planning harmful actions and positioning the attacker as the protector. The attack is decomposed into multi-turns, where the attacker firstly describes others’ harmful intent and seeks prevention, secondly asks about possible evidence items, finally requests an example harmful plan for comparison.
@@
AdvBench [166] and HarmBench [177] are two important benchmarks for jailbreaking LLMs. AdvBench contains a set of 500 harmful behaviors formulated as instructions, as well as a collection of 500 strings that reflect harmful or toxic behavior. HarmBench contains 510 unique harmful behaviors, split into 400 textual behaviors and 110 multimodal behaviors. Both datasets are not intentionally curated for multi-turn jailbreaks, but offer representative jailbreak tasks. Based on AdvBench and HarmBench, Russinovich et al. (2024) [170] manually crafted crescendo multi-turn prompts, and Ren et al. (2024) [171] scales the “attack chain” generation process by employing LLMs via self-talk and six clues from Latour’s actor-network theory. Ren et al. (2024) [171] then compiles the data into SafeMTData. Gibbs et al. (2024) [172] employs word substitution cipher approach [178] to process data from HarmBench with Mixtral-8x7b to isolate the impact of the multi-turn prompting structure, and [173] uses decomposed AdvBench for jailbreaking. In constrast, [174] uses HarmfulQ dataset [179], which comprises 200 explicit harmful questions in English, to process for multi-turn jailbreak. In [175], AdvBench is used for training the jailbreak model and JBB [180] is used for evaluation. In [176], the authors leverage the Beavertails dataset [181], which contains malicious queries across 14 categories designed to test a model’s refusal capability, and employ sentence transformers [182] to generate multi-turn dialogues and harmful actions for GPT-4o.
@@
While researchers have developed numerous multi-turn methodologies to jailbreak LLMs, effective defense strategies specifically designed against such multi-turn attacks remain scarce. Conventional defense methods (e.g., perplexity filters, input/output filters, rephrase/retokenize, rand-drop) demonstrate limited efficacy against these complex, extended conversational interactions. To address this challenge, Yu et al. (2024) [183] shows that system prompts and Chain-of-Thought (CoT) [48] can partially counteract attacks by refusing to answer harmful queries—albeit with lower helpfulness scores—while Gibbs et al. (2024) [172] finds that NeMoGuardrails [184] can be overzealous even with benign prompts, and Liu et al. (2024) [174] notes that stronger defenses often compromise usability. To summarize, robust guardrails against multi-turn jailbreaks remain scarce, and even the limited defenses available often compromise usability by hindering benign interactions. This stark trade-off underscores the urgent need for adaptive, context-aware strategies that can maintain both security and user-friendly performance.
@@ 3 Improvements
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=gray!60, rounded corners=4pt, align=left, minimum width=4em, edge+=purple!40!blue, line width=1.2pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=8.0em,font=,fill=purple!10,, where level=2text width=10.0em,font=,fill=orange!15,, where level=3text width=8.0em,font=,fill=teal!10,, where level=4text width=8.0em,font=,fill=red!10,, [Improvements, ver, fill=green!30, font=, minimum width=12em, minimum height=3em, [Model-Centric
Approaches, align=center, font=, minimum width=11em, minimum height=4em, [ In-Context Learning, text centered [ICL [185], CoT [48], leaf, text width=42em] ] [Fine-tuning, text centered [Vicuna [186], UltraChat [81], PlatoLM [187] , Parrot [24] , [188] [188] , ChatGLM2 [189] ,
CodeSteer [65], leaf, text width=42em] ] [Reinforcement Learning, text centered [Single-Turn RL, text centered [PPO[190], DPO[191], CaPO[24], ACT[192] , leaf, text width=32.5em] ] [Multi-Turn RL, text centered [DMPO[193], M-DPO/M-KTO[52], ArCHer[194], SCoRe[195], REFUEL[196] , leaf, text width=32.5em] ] ] [New Architectures, text centered [CachedTransformer [197], MemBART [198], Transformer-XL [199], RMTransformer [200],
HMT [201], RWKV [202, 203], leaf, text width=42em] ] ], [External Integration
Approaches, align=center, font=, minimum width=11em, minimum height=4em, [Memory Augmentation, text centered [MemPrompt [204], Longmemeval [205], Position [206], MemTree [207], leaf, text width=42em] ] [RAG, text centered [Wizard of Wikipedia [208], [209] [209], [210] [210], BlenderBot 2.0 [211],leaf, text width=42em] ] [Graph Integration, text centered [SinLG [212], [213], DR.KNOWS  [214], SURGE  [215],
Paths-over-Graph [216], GNN-RAG  [217], GNN-RAG  [218], Pseudo-Knowledge Graph [219], leaf, text width=42em] ] ], [Agent-Based
Approaches, align=center, font=, minimum width=11em, minimum height=4em, [Single Agent, text centered [ReAct[220], Toolformer[221], HuggingGPT[222], Reflexion[223], Voyager[224] , leaf, text width=42em] ] [Multi Agents, text centered [CAMEL[225], ChatDev[226], SELF-COLLABORATION[227], MetaGPT[228], Multi-agent Debate[229],
AutoAgents[230], AgentVerse[231] , leaf, text width=42em] ] ] ]
@@ 3.1.1 In-Context Learning
In-context learning (ICL) [185] in multi-turn settings yields nuanced outcomes according to recent empirical studies. While providing exemplars usually aids single-turn tasks, naive in-context prompting with interactive multi-turn examples can sometimes hurt performance in sequential dialog scenarios [29]. For instance, on the AQA-Bench sequential reasoning benchmark, certain models (e.g. LLaMA2 and DeepSeek-LLM in a coin-guessing game environment) actually saw their accuracy drop when given a single in-context demonstration, recovering only as more examples were provided. This surprising inversion of the typical few-shot benefit is attributed to the multi-round nature of the examples, which may cause overfitting to specific interaction trajectories. In other words, multi-turn example formats behave differently from standard one-turn Q&A prompts. Consistent with this, dedicated dialogue benchmarks like MT-Bench [21] and its fine-grained successor MT-Bench-101 [30] highlight that multi-turn interactions demand evaluating facets such as context retention, proactivity, and adaptability that one-shot QA tests often overlook.
@@
Across diverse domains, a variety of prompting techniques have been explored to improve multi-turn LLM interactions, with mixed empirical results. In code generation, interactive multi-turn prompting has proven beneficial: the InterCode [22] benchmark showed that strategies like step-by-step reasoning (e.g. Chain of Thought [48], ReAct) or iterative plan-and-refine prompting yield concrete performance gains on coding tasks by leveraging compiler feedback over turns. In the clinical domain, role-playing a domain expert during training enabled notable gains – Clinical Camel [111], which uses dialogue-based knowledge encoding to infuse medical texts into a Q&A format, surpasses GPT-3.5 on multiple medical benchmarks in few-shot evaluations. By contrast, persona-driven prompts do not universally boost objective accuracy: a systematic study found that adding a persona (e.g. instructing “You are a helpful assistant…”) in system prompts generally failed to improve factual question-answering performance [232]. Nevertheless, persona-centric role-play can enhance conversational quality and engagement in the right context. For example, CharacterChat [78] employs MBTI-aligned personas with preset behaviors and dynamic memory; its role-play prompts facilitated effective personalized support dialogues, demonstrating remarkable efficacy in providing tailored social support. Finally, explicitly steering the dialogue structure has shown promise in certain multi-turn tasks. The StratL [137] approach, which optimizes prompts to guide an LLM tutor along a predefined teaching plan, successfully induced a more pedagogically effective multi-round tutoring strategy (Productive Failure) in practice.
@@
These findings collectively suggest that the impact of ICL and prompt design on multi-turn performance is highly context-dependent – improvements emerge when the prompting strategy is well-aligned with the task’s interactive dynamics, whereas misaligned or naive prompts may even degrade performance in complex multi-turn environments.
@@ 3.1.2 Supervised Fine-Tuning
Classic supervised fine-tuning (SFT) methods like InstructGPT [190] and FLAN [233], along with parameter-efficient techniques such as AdapterDrop [234] and LoRA [235], have driven major progress in single-turn LLM tasks. For a survey on SFT, see [236]. However, to enhance LLMs’ multi-turn interaction ability, SFT must be modified; as Wang et al. (2023) [23] point out, treating each round independently can lead to context forgetting and incoherent responses can lead to context forgetting and incoherent responses in multi-turn dialogues. This limitation has motivated the development of tailored strategies that both leverage realistic multi-turn data and adjust training methods to fully exploit such data.
Moreover, SFT remains one of the most commonly used techniques to boost LLMs’ multi-turn performance. Several studies, as discussed in §2, suggest that incorporating domain-specific multi-turn interaction datasets into SFT can further improve LLM performance. For example, research in multi-turn instruction following (including math reasoning and multi-lingual context) [30, 32, 31], role-play [81, 82] and clinical dialogues [111, 107, 109, 122, 132, 127, 126, 117, 165, 53, 144] demonstrates promising gains. These findings underscore the need for further research to improve fine-tuning efficacy specifically in multi-turn settings.
@@
Recent works have addressed multi-turn challenges by generating datasets that capture natural dialogue flows. For example, Vicuna [186] fine-tunes on user-shared ChatGPT conversations to retain genuine multi-turn interactions, though scaling such data remains costly. Other methods, such as those in UltraChat [81], use self-chat to create extensive multi-turn data; however, these auto-generated dialogues can be overly scripted and lack diversity. To mitigate these issues, approaches like PlatoLM [187] incorporate sophisticated user simulators (e.g., the “Socratic” agent) to produce more dynamic and realistic multi-round dialogues, thereby enhancing topic shifts and follow-up naturalness. In addition, Parrot [24] generates multi-turn data by first mimicking human asking styles for query generation, then constructing negative responses that simulate context neglect or misunderstanding, and finally combining this with context-aware preference signals for fine-tuning.
@@
In parallel with dataset curation, new fine-tuning strategies have been proposed to fully leverage multi-turn data, including aspects of modified loss functions, longer context length, and more efficient training. For example, Vicuna [186] employs a modified loss function and extended context lengths, supported by gradient checkpointing [237] and FlashAttention [238], to handle long dialogue histories efficiently. Similarly, ChatGLM2 [189] extends context length with comparable techniques while also reducing GPU memory costs through multi-query attention [239] and causal masking strategies. More generally, Teng et al. (2024) [188] propose a fine-tuning method that optimizes cross-entropy loss combined with KL divergence across all dialogue turns. Instead of focusing solely on the final turn, this approach leverages the entire conversation, resulting in more coherent outputs and efficient training. Additionally, along with CodeSteer, Chen et al. (2025) [65] address the issue of multi-round gradient cancellation, where gradients from early rounds may cancel out those from later, more informative rounds, by doubling the weights of the final two rounds during fine-tuning. This strategy ensures that the most influential guidance steps drive the model update, improving its ability to select the optimal initial step.

### v5

@@ 2.2.4 Conversational Engagements in Jailbreak
Most prior research has focused on single-turn jailbreaks, in which adversaries use a single prompt—often few-shot—to trigger harmful content. For example, optimization-based methods by Zou et al. (2023) and Liu et al. (2024e) utilize gradient-derived gibberish suffixes to elicit such outputs, but these techniques depend on internal token probability knowledge and are not applicable to closed-source models. To explore alternative strategies, subsequent studies (Carlini et al., 2023; Yu et al., 2023) have investigated unconventional communication patterns, such as role-playing scenarios, and developed multi-turn conversational methods that leverage the entire dialogue history. These approaches demonstrate that multi-turn jailbreaking, where even seemingly innocuous prompts contribute to later interventions, presents a far more complex challenge.
@@
Multi-turn jailbreak can employ more diverse strategies. Crescendo (Russinovich et al., 2025; Ren et al., 2025) is one of the strategies, which implicitly instruct victim LLMs to provide harmful information. The intuition behind this strategy is that LLMs agreeing to a small, initial request increases the likelihood of complying with subsequent, larger demands. By asking innocuous but implicitly suggestive questions, the token distribution of LLMs shifts towards the direction where tokens containing more harmful information are more likely to be generated. The multi-turn interactions as a whole jailbreaks LLMs instead of one specific questions plays the role. Besides, the implicitness of questions and the order to present the questions to LLMs are important to the success of jailbreak. Following the crescendo idea, Ren et al. (2025) generalizes the crescendo method used in Russinovich et al. (2025), where in Russinovich et al. (2025) fixed and human-crafted seed instances are needed to generate attacks, making it challenging to generate diverse attacks. In Ren et al. (2025), the novelty focuses on the process of self-discovering diverse attack clues inside the model’s prior knowledge via network structures and the classification of the clues, by constructing the two-layer relation tree according to Latour’s actor-network theory (Latour, 2005).
@@
Another important strategy for multi-turn jailbreaks is to decompose the one harmful prompt into several pieces where each contains less malicious contents (Gibbs et al., 2024; Zhou et al., 2024c; Liu et al., 2024c). Therefore, language models can incrementally generate harmful content through multi-turn dialogue. By decomposing the original malicious query into several less harmful sub-questions can evade the guardrail of LLMs and induce harmful responses. Due to the in-context learning capabilities, harmful knowledge can be gathered together in the final turn. Using such decomposition, alignment is quite successful for each turn in a multi-turn dialogue, except the final turn when all responses are gathered and combined. However, the cumulative harmful content across the dialogue results in an overall alignment failure. Following the idea, Wang et al. (2024a) train a red-team jailbreak agent through the interactions with target LLMs to generate decomposed yet coherent jailbreak prompts. Jiang et al. (2024b) propose a scenario-based jailbreak method to disguise the malicious intent from guardrails of LLMs, assuming LLMs can only detect direct harmful intent but would be misled if the attacker creates a scenario claiming that others are planning harmful actions and positioning the attacker as the protector. The attack is decomposed into multi-turns, where the attacker firstly describes others’ harmful intent and seeks prevention, secondly asks about possible evidence items, finally requests an example harmful plan for comparison.
@@
AdvBench (Zou et al., 2023) and HarmBench (Mazeika et al., 2024) are two widely used benchmarks for jailbreaking LLMs. AdvBench provides 500 harmful behavioral instructions and 500 harmful/toxic strings, while HarmBench contains 510 unique harmful behaviors spanning 400 textual and 110 multimodal instances. Though neither is designed specifically for multi-turn jailbreaks, both serve as representative testbeds. Building on them, Russinovich et al. (2025) manually craft Crescendo multi-turn prompts, while Ren et al. (2025) scale attack-chain generation via LLM self-talk guided by six clues from Latour’s actor-network theory, compiling the results into SafeMTData. Gibbs et al. (2024) apply a word substitution cipher (Handa et al., 2024) to HarmBench data processed with Mixtral-8x7b to isolate the effect of multi-turn structure, and Zhou et al. (2024c) use a decomposed version of AdvBench for jailbreaking. In contrast, Liu et al. (2024c) draw from HarmfulQ (Shaikh et al., 2023)—200 explicit harmful questions in English—for multi-turn jailbreak construction. Wang et al. (2024a) use AdvBench for training and JBB (Chao et al., 2024) for evaluation, while Jiang et al. (2024b) leverage the BeaverTails dataset (Ji et al., 2023)—covering malicious queries across 14 refusal-test categories—and apply sentence transformers (Ni et al., 2022) to generate multi-turn dialogues and harmful actions targeting GPT-4o.
@@
While numerous multi-turn jailbreak methods have been proposed, effective defenses specifically targeting such attacks remain scarce. Conventional approaches—perplexity filters, input/output filters, rephrase/retokenize, and random token dropping—show limited efficacy against extended conversational attacks. Among targeted efforts, Yu et al. (2024a) show that system prompts combined with Chain-of-Thought reasoning (Wei et al., 2022) can partially counteract attacks by refusing harmful queries, albeit at the cost of reduced helpfulness. Gibbs et al. (2024) find that NeMoGuardrails (Rebedea et al., 2023) can be overzealous, flagging even benign prompts, and Liu et al. (2024c) observe that stronger defenses frequently compromise usability. Collectively, these findings reveal a stark security–usability trade-off, underscoring the urgent need for adaptive, context-aware defense strategies that maintain both robustness and user-friendly performance.
@@ 3.1.1 In-Context Learning
In-context learning (ICL) (Dong et al., 2024a) in multi-turn settings yields nuanced outcomes according to recent empirical studies. While providing exemplars usually aids single-turn tasks, naive in-context prompting with interactive multi-turn examples can sometimes hurt performance in sequential dialog scenarios (Yang et al., 2025b). For instance, on the AQA-Bench sequential reasoning benchmark, certain models (e.g. LLaMA2 and DeepSeek-LLM in a coin-guessing game environment) actually saw their accuracy drop when given a single in-context demonstration, recovering only as more examples were provided. This surprising inversion of the typical few-shot benefit is attributed to the multi-round nature of the examples, which may cause overfitting to specific interaction trajectories. In other words, multi-turn example formats behave differently from standard one-turn Q&A prompts. Consistent with this, dedicated dialogue benchmarks like MT-Bench (Zheng et al., 2023) and its fine-grained successor MT-Bench-101 (Bai et al., 2024) highlight that multi-turn interactions demand evaluating facets such as context retention, proactivity, and adaptability that one-shot QA tests often overlook.
@@
Across diverse domains, a variety of prompting techniques have been explored to improve multi-turn LLM interactions, with mixed empirical results. In code generation, interactive multi-turn prompting proves beneficial: the InterCode benchmark (Yang et al., 2023) shows that step-by-step reasoning strategies such as Chain-of-Thought (Wei et al., 2022) and ReAct, or iterative plan-and-refine prompting, yield concrete performance gains by leveraging compiler feedback across turns. In the clinical domain, Clinical Camel (Toma et al., 2023) uses dialogue-based knowledge encoding to cast medical texts into a Q&A format, surpassing GPT-3.5 on multiple medical benchmarks in few-shot evaluations. By contrast, persona-driven prompts do not universally improve objective accuracy: a systematic study found that adding a persona via system prompts generally fails to boost factual question-answering performance (Zheng et al., 2024a). Nevertheless, persona-centric role-play can enhance conversational quality in the right context—CharacterChat (Tu et al., 2023) employs MBTI-aligned personas with preset behaviors and dynamic memory, facilitating effective personalized social support dialogues. Finally, explicitly steering dialogue structure shows promise: StratL (Puech et al., 2025) optimizes prompts to guide an LLM tutor along a predefined teaching plan, successfully inducing a more pedagogically effective multi-turn tutoring strategy in practice.
@@
Collectively, these findings suggest that the impact of ICL and prompt design on multi-turn performance is highly context-dependent: gains emerge when the prompting strategy aligns well with the task’s interactive dynamics, while misaligned or naive prompts can degrade performance in complex multi-turn settings.
@@ 3.1.2 Supervised Fine-Tuning
Classic supervised fine-tuning (SFT) methods like InstructGPT (Ouyang et al., 2022) and FLAN (Chung et al., 2024), along with parameter-efficient techniques such as AdapterDrop (Rücklé et al., 2021) and LoRA (Hu et al., 2022), have driven major progress in single-turn LLM tasks. A comprehensive survey of SFT is provided by (Zhang et al., 2026). However, to enhance LLMs’ multi-turn interaction ability, SFT must be modified; as Wang et al. (2024f) point out, treating each round independently can lead to context forgetting and incoherent responses can lead to context forgetting and incoherent responses in multi-turn dialogues. This limitation has motivated the development of tailored strategies that both leverage realistic multi-turn data and adjust training methods to fully exploit such data.
SFT remains one of the most widely used techniques for improving multi-turn LLM performance. As discussed in §2, incorporating domain-specific multi-turn interaction data into SFT consistently yields further gains. Evidence spans multi-turn instruction following—including math reasoning and multilingual contexts (Bai et al., 2024; Maheshwary et al., 2025; Liang et al., 2024)—role-play (Ding et al., 2023; Occhipinti et al., 2024), and clinical dialogue (Toma et al., 2023; Zhang et al., 2023; Bao et al., 2023; Yang et al., 2024; Liao et al., 2023; Ye et al., 2023; Zhao et al., 2024; Pieri et al., 2024; Wang et al., 2024d; Macina et al., 2023; Scarlatos et al., 2025). These results underscore the need for further research into fine-tuning strategies tailored specifically to multi-turn settings.
@@
Recent works have addressed multi-turn challenges by generating datasets that capture natural dialogue flows. For example, Vicuna (Chiang et al., 2023) fine-tunes on user-shared ChatGPT conversations to retain genuine multi-turn interactions, though scaling such data remains costly. Other methods, such as those in UltraChat (Ding et al., 2023), use self-chat to create extensive multi-turn data; however, these auto-generated dialogues can be overly scripted and lack diversity. To mitigate these issues, approaches like PlatoLM (Kong et al., 2024b) incorporate sophisticated user simulators (e.g., the “Socratic” agent) to produce more dynamic and realistic multi-round dialogues, thereby enhancing topic shifts and follow-up naturalness. In addition, Parrot (Sun et al., 2024) generates multi-turn data by first mimicking human asking styles for query generation, then constructing negative responses that simulate context neglect or misunderstanding, and finally combining this with context-aware preference signals for fine-tuning.
@@
Alongside dataset curation, new fine-tuning strategies have been proposed to fully exploit multi-turn data, targeting loss design, context length, and training efficiency. Vicuna (Chiang et al., 2023) employs a modified loss function and extended context lengths, supported by gradient checkpointing (Chen et al., 2016) and FlashAttention (Dao et al., 2022) to handle long dialogue histories efficiently. ChatGLM2 (GLM et al., 2024) similarly extends context length while reducing GPU memory costs via multi-query attention (Shazeer, 2019) and causal masking. More generally, Teng et al. (2024) optimize a combined cross-entropy and KL divergence loss across all dialogue turns rather than the final turn alone, yielding more coherent outputs and more efficient training. Finally, Chen et al. (2025d) address multi-round gradient cancellation—where gradients from early turns can cancel those from later, more informative ones—by doubling the loss weights of the final two rounds, ensuring that the most influential guidance steps drive model updates and improving selection of optimal initial actions.